On December 31 of Year 1, a company reports accounts receivable of $110,000. During Year 2, the company has net credit sales of $1 million. On December 31 of Year 2, the company reports accounts receivable of $140,000.
  
What is this company’s accounts receivable turnover ratio for Year 2?

A. 7.14

B. 8.00

C. 8.33

D. 9.09
Answer #1
Accounts receivable turnover = Net sales/Average receivables
Average receivables = (110000+140000)/2 = 125000
Accounts receivable turnover = 1000000/125000 = 8
Answer: Option [b]